# Pre-Columbian Settlements Shaped Palm Clusters in the Sierra Nevada de Santa Marta, Colombia

### v1

@@ Introduction
Study Area
@@
We selected a study area of approximately 765 Km2 in Colombia, which is known to contain Tairona occupation. We collected archaeological legacy data, harmonizing geospatial location of historical maps and archaeological reports to estimate areas with possible archaeological sites with low and medium certainty [28]. High-certainty locations were identified through pedestrian surveys, as detailed in [50, 28]. Figure 1 presents the study area at various zoom levels for better spatial context.
@@
Within this region, we used high-resolution satellite imagery from the Pléiades Neo constellation acquired between 2022 and 2024 to annotate the locations of star-shaped palm trees. We utilized the panchromatic band, which offers the highest spatial resolution (30 cm), making it particularly well-suited for detecting individual palm trees. A total of 37 scenes covering the study area were downloaded and used for analysis. A 69.5 Km2 subset was annotated with the locations of star-shaped palm trees and used to train and test the object detection model. Because detection was at the canopy layer (top forest layer) and the study area is a montane forest, the most likely palm detected is Dictyocaryum sp. This genus is known for its tall, solitary growth form and prominent canopy presence in Andean cloud forests [37, 17], making it more evident from understory palms like Bactris sp. or lowland species such as Phytelephas sp. To streamline the annotation process and minimize the risk of missing any palm trees, the area was divided into 200x200 m grid. Human annotators systematically scanned each grid and marked individual palm trees bounding boxes. In total, 24,430 palm trees were annotated. (Figure 1, bottom-center).
@@
First, we trained a YOLOv10x-based deep learning model [52] on 30 cm resolution Pléiades Neo satellite imagery to detect palm trees (Figure 2). YOLOv10x, an object detection architecture, delivers high accuracy for small and densely clustered classes like the palm trees in this study. The labeled palm area was divided into two spatially distinct regions (Figure 2). The lower region, comprising 47% of the annotations, was used for training and validation, while the upper region, containing the remaining 53%, was reserved for testing. This spatial separation aimed to reduce the risk of spatial bias during classification. Each region was further subdivided into 640×640 pixel patches for model input. Within the training and validation region, 95% of the patches were used for training and the remaining 5% were used for validation to guide hyperparameter tuning and model selection. The best-performing model was applied to the test patches for evaluation, as well as to the full study area, for palm tree detection.
@@
Next, we applied Hierarchical Density-Based Spatial Clustering of Applications with Noise (HDBSCAN) [3] to identify spatial clusters of palm trees using 100 as minimum number of points required to form a cluster. This algorithm handles clusters of varying densities, it does not require a parameter for neighborhood radius, and it is robust to noise and outliers. We linked these clusters to past human activity by associating them with the centroids of known archaeological sites and low-location-certainty archaeological zones within the convex hull boundaries of each palm cluster.
@@
To compare palm concentrations, we generated centroids for archaeological sites and low-location-certainty archaeological zones. We also generated random centroids in the largest palm cluster and in areas showing clustered palms without reported archaeological sites. Palm concentration differences were measured using a Inverse Distance Weighting (IDW) score, which models the density of nearby palm trees by assigning higher weights to those located closer to each centroid. Palm trees within a 1000 m radius of each centroid were included in the IDW calculations, ensuring that the analysis captured localized patterns while excluding influence from distant and ecologically distinct areas. This radius is an conservative estimate of human mobility based on recently reported walking distances and travel times between Indigenous settlements (see introduction). The score is calculated as $G=\sum_{i}\frac{1}{d_{i}^{w}}$, where $G$ is the IDW score, $d_{i}$ is the distance from the selected point to palm tree $i$ (only if $d_{i}\leq 1000$ meters), and $w$ is a decay factor that gives more weight to closer trees. We use $w=1$ in our analysis. Finally, we used bootstrapping techniques to compare palm concentrations around archaeological centroids with those in control zones—palm areas without known archaeological features. For this, we generated 100 control points, randomly drawn from palm cluster regions without known archaeological features. We also generated an additional set of 100 control points within the largest palm cluster but outside 1 km buffers around the archaeological sites. For both control sets and the archaeological centroids within the largest palm cluster, we created 1,000 bootstrapped datasets. Each dataset consisted of 17 observations drawn with replacement, matching the number of archaeological centroids in the largest cluster.
@@
We evaluated the consistency of the palm detection model by comparing the elevation of detected palms with the elevation distribution of Dictyocaryum sp. occurrences recorded in the Global Biodiversity Information Facility (GBIF) database [12]. Elevation data for both sets were extracted from the SRTM dataset at 1 arc second ($\sim$30 meter) resolution [34]. An IQR-based outlier filtering on the SRTM elevation was applied to the GBIF baseline. GBIF records served as proxy for the species’ natural elevational range, likely reflecting areas less affected by pre-Columbian human activity. We then compared GBIF data with the elevations of palm trees detected by our deep learning model within a 1000 m radius of archaeological centroids in the largest palm cluster. We also examined the elevation data for all palm trees located in areas showing clustered and non-clustered palms without archaeological sites.
@@
To assess the potential relationship between the number of modern human activities and the number of palm trees detected within the study area, we conducted a spearman rank correlation between the number of modern buildings detected in a 200x200 m cell grid and the detected palms (Figure 3B). The analysis used a dataset containing automatic detected building from satellite imagery between 2014 and 2024 [32] and palm trees detected per spatial unit (Figure 3B).
Figure 2: PABAM methodology, which incorporates the YOLOv10 architecture. After selecting the best-performing model based on validation data, we applied it to the test set for performance evaluation, as well as to the entire study area for palm tree detection. Detected palm trees were clustered using the HDBSCAN algorithm, followed by an IDW analysis to compare palm tree densities. Each cluster with its corresponding cluster ID is shown on the bottom-left frame. Left panels include material © CNES (2024), Distribution Airbus DS. Right panel includes Sentinel-2 cloudless - https://s2maps.eu by EOX IT Services GmbH (Contains modified Copernicus Sentinel data 2016 & 2017).
Results
Palm Tree Detection Model Performance
We evaluated the performance of the model using different confidence thresholds. Precision and recall metrics are reported in Figure 2. In our application, minimizing false positives was more important than maximizing recall because high false positive rates in palm tree detection could produce artificial clusters and misleading density estimates. Since clusters should reflect true patterns with high precision, and our goal was also to develop a generalizable model for diverse geographic regions, we prioritized precision over recall. Therefore, a confidence threshold of $\geq 0.4$ was selected to reduce false positives. The final palm tree bounding box predictions for the entire study area, based on the $\geq 0.4$ threshold, are shown in Figure 2.
Analysis of Palm Distribution and Archaeological Sites
To evaluate the relationship between palm tree distribution and past human presence, we conducted a spatial clustering analysis of palm trees across the study area. Clusters 1, 2, and 15 (located in the southwest, west, and east portions of the bottom-left frame in Figure 2) exhibited spatial overlap with known archaeological features. Because the areas in the western portion of the study area have undergone more modern human intervention, we chose to focus our analysis in the less-disturbed eastern part of the area of study.
Notably, the largest archaeological site (Teyuna), marked by significant infrastructure investment, is located within the largest cluster of palms (Cluster 15), spanning approximately 100 Km2 across both well-lit ridges and steep shaded slopes in the eastern part of the study area (Figure 3C). This cluster also contains 17 archaeological zones recorded as points or polygons. We calculated the centroids of these sites to serve as reference points for spatial analysis. This resulted in 17 reference points used in our study. Archaeological sites within the largest palm cluster are located in areas with significantly higher IDW values than random locations within palm clusters lacking archaeological features (Figure 3A). To estimate the extent of ancient human management influence in Cluster 15, we compared palm tree densities at 100 randomly selected points within the cluster but outside a 1 km buffer around archaeological sites with densities at the 17 archaeological points in the same cluster. The confidence intervals overlapped substantially, indicating that the difference in palm density is not statistically significant (Figure 3A).
Extended Analysis Across Regional Archaeological Centroids
Beyond the localized analysis of archaeological centroids near Teyuna, we expanded the IDW analysis to include four additional archaeological centroids situated in geographically distinct regions of the study area. These centroids were spatially independent from both the Teyuna-associated centroids and the control points used in the initial analysis. One of these centroids, in the southwest of the study area corresponds to El Congo-Ciudad Antigua [43], a previously known archaeological site. The remaining three centroids were estimated based on nearby archaeological sites reported in legacy data in recorded in our surveys. The IDW values for the centroids in the west, southwest, and south regions were 0.04, 0.14, and 0.05 respectively. In contrast, the Teyuna centroid yielded a IDW value of 20.4. These results suggested that palm tree density around Teyuna is substantially greater than in the other regions.
Figure 3: Human activity and palm distribution. (A) Confidence intervals for bootstrapped means (Bootstrap samples = 1000; sample size = 17). The subsets shown are IDW means for archaeological centroids inside Cluster 15 (left), 100 random points in Cluster 15 outside 1 km archaeological buffers (center), and 100 random points in clusters without overlap with archaeological features (right). Bootstrapped means are shown as text labels to the right of each bullet graph. (B) Spearman correlation between modern buildings [32] and palm detections in each 200x200 m grid inside the study area. (C) Elevation distributions across subsets of detected palms and the GBIF baseline for the distribution of Dictyocaryum in South America. Violin plots (left) display Gaussian kernel density estimates for each subset, with overlaid box plots indicating the median (white line) and mean (red dot). (D) The heatmap presents pairwise Dunn test results between subsets following a significant Kruskal-Wallis test.
Correlation Between Modern Buildings and Palm Tree Counts
The Spearman correlation between the number of modern human buildings and the number of palm trees detected in each 200x200 m cell grid and the detected palms (Figure 3B) was low ($p<0.001$), suggesting little to no relationship between the presence of modern buildings and the distribution of palm trees in the study area.
Elevation Analysis
The violin plots in Figure 3C show that palm trees located within a 1 km radius of archaeological sites in Cluster 15 tend to occur at lower elevations compared to those outside this buffer zone within the largest palm cluster in the study area. For reference, the centroids of archaeological locations within Cluster 15 span elevations ranging from 819 to 1570 m, with a mean elevation of 1113 m. Furthermore, both elevation distributions—palms near archaeological sites and those outside the buffer—show more frequently values at lower than those of other palm clusters without archaeological overlap, as well as non-clustered palm trees in the study area. All elevation distributions of the detected palm subsets fall within the range of the GBIF baseline, suggesting that the palm detection is consistent capturing the overall variability of palm distribution. In addition, all median elevation values of clustered palms that overlap with archaeological sites are below the GBIF baseline median. In contrast, only the clusters of palms without archaeological overlap exhibit median elevation values higher than that of the GBIF baseline.
Discussion
Previous studies typically examined broad geographical regions using low-resolution data [26, 44], which can miss finer-scale patterns of human–environment interaction. Our high-resolution analysis across a mid-sized area reveals a spatial correlation between the pre-Columbian populations and the densest cluster of palm trees in the Neotropic forest of the SNSM. This correlation is not consistent across all sites but is particularly pronounced around Teyuna, where pre-Columbian populations invested significantly in infrastructure development. Settlement data from the southwestern part of our study area indicate that most archaeological sites are relatively small ($\leq$ 1 ha), with a maximum surface of around 5 ha [50, 47]. This surface is considerably smaller than the main site of Teyuna ($\sim 33$ ha). The area around Teyuna contains a significantly higher number of individuals from the detected plant taxa than would be expected under a random distribution model, suggesting that pre-Columbian activities shaped this pattern. While Teyuna is occupied from around 400 CE [13], paleoecological reconstructions show an increase in palm pollen [49, 19], coinciding with the first centuries (1100 and 1250 CE) of the Tairona period. Given that the estimated ancient human management area at the Teyuna site alone is around 100 ha [43], incorporating the large surrounding cluster of detected palms suggests a much broader area of influence. This palm cluster spans approximately 100 Km2 suggesting that the extent of activity around Teyuna, including paths, cultivation, resource collection, and general forest management could be up to two orders of magnitude larger than what archaeological evidence alone indicates. This finding complements the temporal data from the palynological analysis and is aligned with the period of major infrastructural investment in the SNSM. Similar dynamics have been observed in many parts of the Amazon, where archaeological earthworks coincide with the hyperdominance of useful species [39, 26], and in the highlands of Colombia, where activities around pre-Hispanic population centers likely impacted high-altitude lakes [51].
Teyuna and adjacent sites located within the largest identified palm cluster exhibit significantly higher palm densities compared to randomly selected points inside clusters without evidence of prehispanic occupation within the same region. This pattern suggests that pre-Columbian activities that managed the forest, and possibly opened the canopy, favored these palms. Along with canopy openness, the typically short dispersal distances in palms combined with the low density of other species probably contributed to the higher local density. In the case of Dictyocaryum lamarckianum, its stilt roots provide an increased mechanical support on steep hillsides, again offering an advantage to occupy the cleared landscape. Following initial settlement and palm proliferation, the technological utility of palms likely enhanced the value of the area and influenced subsequent settlement decisions by offering valuable resources—such as materials for construction, containers, thatch, and edible parts like palm hearts [37, 33]—making these areas more suitable for long-term habitation and infrastructure investment.
Building on the extended analysis across multiple archaeological sites, and given Teyuna’s historical prominence as a major pre-Columbian center [43, 13], it is plausible that anthropogenic influences—such as the planting, cultivation, or preferential retention of palms—were more intensive in this region. These practices may have constituted a form of low-intensity management or stewardship of plant communities [5]. Our results suggest that the legacy effects of pre-Columbian human activity on palm distribution are not uniform across the landscape. Instead, they appear to be spatially variable, with a more pronounced presence persisting near major centers such as Teyuna.
Elevation analysis also supports the idea that prehistoric human activity contributed to localized ecological modifications that promoted palm persistence or proliferation at lower altitudes than the baseline data. The presence of Pre-Columbian Indigenous populations possibly facilitated the creation of micro-environmental conditions, such as access to light, modified hydrology, reduced interspecific competition, and/or direct management practices, that favored the survival and expansion of clusters of Dictyocaryum palms at lower altitude sites. This presence of palms supports the hypothesis that pre-Columbian people management of the forest has left a durable ecological imprint. Although such occurrences are unlikely to act as standalone predictors for the presence of archaeological sites, they offer valuable supplementary information when integrated with broader environmental and multispectral and hyperspectral datasets. The palm distribution detected in our area of work of the SNSM is an example of long-term human-ecological interactions and niche construction [22], where sometimes subtle but persistent anthropogenic influences shape the distribution of key plant species in time and space.
The current palm distribution is the result of ecological processes that unfolded during the Pre-Columbian occupation of these sites. The composition of the SNSM forest in areas without reported human presence indicates that palms are a key component of the floristic assemblage but their frequency in the vegetation is relatively low [11, 7]. Although short-term successional dynamics following forest disturbance are relatively well understood for lowland Amazonia [35]—for instance, the dominance of early successional species depending on disturbance scale—such processes in Neotropical montane environments remain less studied. Our data suggest that patterns of certain plant frequencies that coincide with and/or nearby archaeological sites are most likely the result of human agency rather than forest regeneration dynamics. More research using plant proxies for plants such as phytoliths extracted from archaeological deposits and soils from nearby areas is needed to understand these dynamics.
Potential applications
Our approach could be extended to analyze other Neotropical forests, such as the Amazon, as well as more arid landscapes with sparse vegetation—where plant species may serve as especially clear ecological markers of past human activity. These markers may reflect lasting ecological legacy of indigenous land use [38] or post-abandonment conditions shaped by earlier human presence [45]. Combining automatic species detection and botanical field surveys with archaeological and paleoenvironmental data can enhance our understanding of how pre-Columbian indigenous settlements in the Neotropics, and contextualize them within interconnected local and global dynamics of past human–environment relationships [16]. Beyond archaeological applications, the model have potentials for efficient palm tree distribution mapping [10] for bio-economic development, land-use planning, and ecological monitoring. Our method, using satellite imagery and deep learning, provides a scalable approach to detect such patterns on wider surfaces and more remote regions. By automatically mapping the distribution and dominance of specific plant species—particularly those with cultural or ecological ties to past human groups—this approach can aid in identifying affiliation of current indigenous groups to land, complementing traditional field-based evidences.
Limitations and future work directions
Although the palm detection model used in this study successfully captured the overall distribution and density patterns of palm trees throughout the canopy of the study area, it occasionally failed to accurately detect individual trees. These errors are primarily caused by visual similarities between palm trees and certain human-made structures, such as crossroads or buildings, that can exhibit textures or shapes resembling the star-like appearance of a palm tree crown.These misclassifications mainly occur in heavily modified landscapes; in well-preserved forests—where this approach is intended—the risk of confusing natural features with roads or buildings is minimal. To address this issue, future work can focus on integrating semi-supervised learning into the detection framework to enhance palm tree detection performance. By combining limited labeled data with a larger corpus of unlabeled imagery, these methods can help models learn more robust and generalizable features. Techniques such as pseudo-labeling [55] and consistency regularization [54] may be particularly effective in leveraging the variability present in unlabeled data, including differences in appearance, lighting, and environmental conditions. This strategy is expected to improve detection accuracy and adaptability to unseen regions and palm tree varieties.
The computer vision model detected a specific crown pattern at the top of the canopy. Palms with different crown patterns or located lower in the canopy were less detected. Some areas in the study region—particularly in the southwest—show signs of modern human intervention, while the area around Teyuna appears more preserved. To further enhance the efficiency and accuracy of palm detection, future research could focus on integrating active learning into the training pipeline [42]. This would involve iteratively training a model on a small labeled set, using the model to identify the most informative unlabeled samples (e.g., those with high prediction uncertainty), acquiring annotations for these selected samples, and retraining the model with the expanded labeled set. This iterative process has the potential to significantly improve performance while minimizing annotation costs.
Conclusion
This study demonstrates an association between palm distribution and archaeological sites in the Sierra Nevada de Santa Marta, Colombia. Using our scalable methodology for identifying potential archaeological features through their ecological signatures we find that palm density is indicative of past human activity, suggesting that a synergetic relation between initial settlement and subsequent palm proliferation influenced decisions on where to continue developing settlements. The significantly higher palm densities around archaeological centroids—particularly at Teyuna—and the observation that site-associated palms occurred at lower elevations than both regional baselines and non-site-associated palms, indicate that pre-Columbian populations actively shaped environments in ways that promoted palm proliferation in areas where plant species competition should be intense. This in turn, may have lowered the logistical costs of establishing infrastructure-heavy settlements in otherwise less accessible locations.

### v2

@@ Conclusion
Pre-Columbian Settlements and Forest Management in the Sierra Nevada de Santa Marta
Pre-Columbian communities of the Tairona cultural tradition transformed the SNSM landscape. Human occupation, dated from 200 to 1600 CE  (?, ?), is divided into two archaeological periods: Neguanje (200–1100 CE) and Tairona (1100–1600 CE). Settlements occur on ridgelines, valley slopes  (?, ?, ?), riverbanks, bays, inlets, flat coastal areas, and deeper canyons  (?, ?). Long-distance mobility between sites within the SNSM has been documented both historically  (?) and in the present, with a 60 km round-trip between two Indigenous communities completed in 18 hours  (?). Entire valleys, typically 30 km long, can be traversed in three days from sea level to the paramo ($\sim 3000$ m elevation) (?). Spanning from coastal lowlands to 2000 m elevation, the Tairona sites feature terraced platforms with curved stone walls, rammed earth foundations, rainwater drainage systems, and circular stone floors for wood-and-thatch structures connected by paved paths  (?, ?, ?, ?, ?, ?). All this represent a significant infrastructure investment given the rugged terrain and the logistical challenges of transporting large amount of stones.
Small settlements (<<3 ha) are scattered across the landscape  (?). Larger sites ($\geq$3 ha) are less common but present in both the Neguanje and Tairona periods. However, estimating the size of settlements in the Neguanje phase is challenging due to the superimposing Tairona structures that cover the earlier deposits  (?). The largest known settlement is Teyuna (Ciudad Perdida or Buritaca-200), occupied in the Neguanje period from around 400 CE  (?). During the Tairona period, it expanded to a settlement of 33 ha with typical regional architecture  (?, ?, ?). A remote sensing analysis estimates the footprint of Teyuna at approximately 100 ha (Figure 1), supporting the role of the settlement as a regional center within an extensive and interconnected system, but also suggesting a broader anthropogenic landscape surrounding the site. This aligns with the models of “low-density agricultural urbanism” seen in Mesoamerica and Southeast Asia  (?) and with evidence from other Neotropical forests such as the Amazon  (?).
Palynological analyses  (?, ?) in a radius of 6 km from Teyuna, along with phytolith studies at the site  (?), indicate that landscape transformation began between 600 and 900 CE, likely for small-scale cultivation of maize (Zea mays) and manioc (Manihot sp.). Phytoliths from Bactris sp. and other palms (Arecaceae) were identified in Teyuna, suggesting the exploitation of such plant resource  (?). Also, an increase in palm pollen between 1100 and 1250 CE in the lowlands  (?), and at 900-1300 m elevations  (?), coincided with the Tairona occupation period. Among palms, Dictyocaryum lamarckianum—common in the SNSM at elevations of 900-2000 m—is used for construction and may represent one of the sources of phytoliths within the recovered assemblages  (?, ?). Palynological evidence also suggests rapid forest regeneration after site abandonment in the 16th century, implying that Tairona settlements had an ecological footprint on modern forest composition  (?).
@@
We selected a study area of approximately 765 km2 in Colombia, which is known to contain Tairona occupation. We collected archaeological legacy data (Suplementary Material, Data S1 and S2), harmonizing geospatial location of historical maps and archaeological reports to estimate areas with possible archaeological sites with low and medium certainty  (?). High-certainty locations were identified through systematic pedestrian surveys, as detailed in previous studies  (?, ?). Figure 1 presents the study area at various zoom levels for better spatial context.
@@
Within this region, we used high spatial resolution satellite imagery from the Pléiades Neo constellation acquired between 2022 and 2024 to annotate the locations of star-shaped palm trees. We utilized the panchromatic band, which offers the highest spatial resolution (30 cm), making it particularly well-suited for detecting individual palm trees. A total of 37 scenes covering the study area were used for analysis. A 69.5 km2 non-contiguous subset was annotated with the locations of star-shaped palm trees and used to train and test the object detection model (Suplemmentary Material, Data S3 and S4). To streamline the annotation process and minimize the risk of missing any palm trees, the area was divided into 200x200 m grid. Human annotators systematically scanned each grid and marked individual palm trees bounding boxes. In total, 24,430 palm trees were annotated. (Figure 1, bottom-center). Since orbital optical imagery is typically acquired at nadir, detection is biased toward the canopy layer. In the montane forest of the study area, the most likely palm detected under these conditions is Dictyocaryum sp. This genus is known for its tall, solitary growth form and prominent canopy presence in Andean cloud forests  (?, ?). This makes it more readily distinguishable than understory palms such as Bactris sp. or lowland species such as Phytelephas sp.
@@
First, we trained the extra-large variant of the “You Only Look Once” (YOLO) object detection architecture, a family of deep learning models designed for fast and accurate object localization and classification in images  (?). Within it’s 10th variant, YOLO introduces several improvements in efficiency and accuracy, and the x configuration represents the most accurate variant, optimized for detecting small or complex objects. We applied this model (YOLOv10x) to 30 cm resolution Pléiades Neo satellite imagery to identify palm trees (Figure 2), taking advantage of its ability to capture fine-scale canopy features. The labeled palm area was divided into two spatially distinct regions (Figure 2). The lower region, comprising 47% of the annotations, was used for training and validation, while the upper region, containing the remaining 53%, was reserved for testing. This spatial separation aimed to reduce the risk of spatial bias during classification. Each region was further subdivided into 640×640 pixel patches for model input. Within the training and validation region, 95% of the patches were used for training and the remaining 5% were used for validation to guide hyperparameter tuning and model selection. The best-performing model based on precision and recall scores was applied to the test patches for evaluation, as well as to the full study area, for palm tree detection (Suplementary Material, Data S5).
@@
Next, we applied Hierarchical Density-Based Spatial Clustering of Applications with Noise (HDBSCAN)  (?) to identify spatial clusters of palm trees using 100 as minimum number of points required to form a cluster. This algorithm handles clusters of varying densities, it does not require a parameter for neighborhood radius, and it is robust to noise and outliers. We linked these clusters to past human activity by associating them with the centroids of known archaeological sites and low-location-certainty archaeological zones within the convex hull boundaries of each palm cluster.
@@
To compare palm concentrations, we generated centroids for archaeological sites and low-location-certainty archaeological zones. We also generated random centroids in the largest palm cluster and in areas showing clustered palms without reported archaeological sites. Palm concentration differences were measured using a Inverse Distance Weighting (IDW) score, which models the density of nearby palm trees by assigning higher weights to those located closer to each centroid. Palm trees within a 1000 m radius of each centroid were included in the IDW calculations, ensuring that the analysis captured localized patterns while excluding influence from distant and ecologically distinct areas. This radius is an conservative estimate of human mobility based on recently reported walking distances and travel times between Indigenous settlements  (?, ?). The score is calculated as $G=\sum_{i}\frac{1}{d_{i}^{w}}$, where $G$ is the IDW score, $d_{i}$ is the distance from the selected point to palm tree $i$ (only if $d_{i}\leq 1000$ meters), and $w$ is a decay factor that gives more weight to closer trees. We use $w=1$ in our analysis. Finally, we used bootstrapping techniques to compare palm concentrations around archaeological centroids with those in control zones—palm areas without known archaeological features. For this, we generated 100 control points, randomly sampled from palm cluster regions without known archaeological features. We also generated an additional set of 100 control points within the largest palm cluster but outside 1 km buffers around the archaeological sites. For both control sets and the archaeological centroids within the largest palm cluster, we created 1000 bootstrapped datasets. Each dataset consisted of 17 observations sampled with replacement, matching the number of archaeological centroids in the largest cluster.
@@
We further evaluated the consistency of the palm detection model by comparing the elevation of detected palms with the elevation distribution of Dictyocaryum sp. occurrences recorded in the Global Biodiversity Information Facility (GBIF) database  (?). Elevation data for both sets were extracted from the Shuttle Radar Topography Mission (SRTM) dataset at 1 arc second ($\sim$30 m) resolution  (?). An interquartile range-based outlier filtering on the SRTM elevation was applied to the GBIF baseline. GBIF records served as proxy for the species’ natural elevational range, likely reflecting areas less affected by pre-Columbian human activity. We then compared GBIF data with the elevations of palm trees detected by our deep learning model within a 1000 m radius of archaeological centroids in the largest palm cluster. We also examined the elevation data for all palm trees located in areas showing clustered and non-clustered palms without archaeological sites.
@@
To assess the potential relationship between the number of modern human activities and the number of palm trees detected within the study area, we conducted a Spearman rank correlation between the number of modern buildings detected in a 200x200 m cell grid and the detected palms (Figure 3B). The analysis used a dataset containing automatic detected building from satellite imagery between 2014 and 2024  (?) and palm trees detected per spatial unit (Figure 3B).